Upgrade to macoS Mojave gnments Pre-material Assignment Pre-material Assignment 《Prev Next > Question 20 Use the 95% ruleand the fact that the summary statistics come from a distribution that is symmetric and bell shaped to find an interval that is expected t0 contain about95% of the data values. A bell-shaped distribution with mean 160 and standard devlatlion 21 The interval is eTextbook and Media Attempts: unlimited Check Arower Send to Gradebook Prev Next
Answer #1

The interval that expected to contains about 95% of the data values is,

mean - 2 * standard deviation = 160 -2*21 = 160- 42 =118

mean + 2* standard deviation =160 +2*21 =160 +42=202

The interval is 118 to 202